OpenStudy (jiteshmeghwal9):

If f(x) be a twice differentiable function & f(0)=0,f"(0)=5, then \[\lim_{x \rightarrow 0}\frac{f(x)-4f(2x)+f(7x)}{x^2}=?\]

OpenStudy (priyar):

again use L-hospital's rule

OpenStudy (priyar):

differentiate the numerator and denominator separately and plug in the given values..

OpenStudy (priyar):

did u understand ? @jiteshmeghwal9

OpenStudy (samigupta8):

Gud explanation priyar....she is ryt....jiteshmeghwal follow d same procedure as described by her

OpenStudy (jiteshmeghwal9):

\[\lim_{x \rightarrow 0}\frac{f'(x)-4f'(2x)+f'(7x)}{2x}\]Is this the first step?

OpenStudy (priyar):

almost..u forgot to apply the chain rule.. for f(2x) and f(7x)..

OpenStudy (jiteshmeghwal9):

plz tell me how to apply chain rule

OpenStudy (samigupta8):

f'x-4f'(2x)*2+f'(7x)*7 is the numerator here

OpenStudy (priyar):

f(2x) differentiation is 2f'(2x) cos we must also differentiate the "2x" inside..ok?

OpenStudy (jiteshmeghwal9):

\[\lim_{x \rightarrow 0}\frac{f(x)-8f'(2x)+7f'(7x)}{2x}\]

OpenStudy (priyar):

good job!

OpenStudy (priyar):

but u forgot to put f'(x)

OpenStudy (priyar):

its ok..differentiate again

OpenStudy (jiteshmeghwal9):

sorry \[\lim_{x \rightarrow 0}\frac{f"(x)-16f"(2x)+49f"(7x)}{2}\]

OpenStudy (priyar):

correct! now put the values given..

OpenStudy (jiteshmeghwal9):

\[{5-80+245 \over 2}={170\over2}={85}\] i gt the answer thx a lot :)

OpenStudy (priyar):

great job!!

OpenStudy (priyar):

u r welcome!